# Self-supervised Learning of LiDAR 3D Point Clouds via 2D-3D Neural Calibration

### v3

@@ IV-D Panoptic Segmentation
To assess the instance-level features learned by our method, we further fine-tune our pre-trained models for the task of panoptic segmentation.
Network Architectures. In this part, we choose the MINKUnet-34 [40] as the 3D point-wise feature backbone. For the fine-tuning process, semantic and instance heads are integrated with the pre-trained 3D backbone, followed by clustering post-processing to identify the instances [10]. This clustering stage leverages semantic predictions to exclude background points, focusing solely on foreground instances such as vehicles, pedestrians, and cyclists. Once the irrelevant points are removed, the remaining points undergo a clustering process to discern distinct instances based on the features from the instance head. For the clustering phase, we opt for the mean shift algorithm [56] with a set bandwidth of 1.2 and a minimum cluster size requirement of 50 points. The other architectures and settings are the same as Section IV-B during pre-training.
Protocol for Fine-tuning on Downstream Tasks. Our methodology employs the AdamW optimizer, utilizing a max learning rate set at 0.2. In the context of panoptic segmentation applied to the SemanticKITTI dataset, the fine-tuning of our model is conducted across various subsets of the annotated training data, specifically 0.1%, 1%, 10%, 50%, and the full 100% subset. For the nuScenes dataset, we fine-tune using both the complete training set and the mini training subset that is provided. The performance of the fine-tuned model is then assessed on the complete validation set of all datasets. We fine-tune the model for fixed 50 epochs.
Results of Fine-tuning. We pre-train the backbone on the SemanticKITTI dataset and fine-tune it on different percentage subsets of SemanticKITTI dataset. As shown in Table VI, our method is consistently better than previous self-supervised pre-training approaches. When the segmentation model is trained with fewer labels, we can observe an obvious improvement compared to the baseline of scratch training.
Generalization of Learned Features. Our study also investigates the adaptability of learned features focusing on panoptic segmentation tasks. We initially pre-train the 3D backbone on the SemantiKITTI dataset and subsequently fine-tune it on both the full and mini training sets of nuScenes. As presented in Table VII, while all approaches enhance performance on the nuScenes dataset, the features generated through our technique demonstrate greater adaptability when employed in a different domain dataset. Furthermore, in comparison to supervised pre-training on SemanticKITTI, our approach exhibits improved results. This finding underscores the enhanced effectiveness and potential of our method in scenarios of transfer learning, surpassing the conventional supervised pre-training methods.
@@ IV-E Ablation Studies
In our ablation studies on the nuScenes dataset, we adhere to the evaluation protocol from ALSO [15]. This involves splitting the nuScenes training set into two parts: ablation-train and ablation-val. Fine-tuning is performed with just 1% of annotations from the ablation-train set. This ensures parameter tuning does not involve the validation set, reserved for comparing against other methods. We limit the training period for these studies to 100 epochs.
@@
Effect of Key Components. We analyze the impact of each component incorporated into our method in Table VIII,. We can observe that our proposed key components benefit the pre-training framework in learning useful representation and yield better performance in downstream tasks. For example, the introduction of learnable transformation alignment increases the mIoU from 35.90% to 36.52%. This indicates that allowing the model to adaptively transform features enhances its ability to capture the inherently non-linear and complex relationship between point cloud data and RGB images. Besides, the supervision of advanced tasks, including overlapping area detection and pose estimation also prompts the network to fully understand the two modalities and learn useful representation. Then we analyze the effect of learnable transformation alignment in detail.
Method | Similarity |
w/o LTA | 0.55 0.17 |
w/ LTA | 0.69 0.13 |
@@
We also provide the quantitative comparison of similarities between features of two modalities in Table IX. Specifically, we compute the cosine similarity between the features of the points and corresponding pixels with and without linear feature transformation, respectively. We can observe that directly computing the cosine similarities yields smaller values for the corresponding points and pixels. However, applying the learnable feature transformation bridges the gap between the two modalities and brings more reasonable similarities.
@@ IV-F Feature Matching Accuracy
We present a comparative analysis of feature matching accuracy using various feature alignment and point-to-pixel matching methods. The mean and standard deviation of error distances are detailed in Table X. To assess the precision of different methods, we calculate the matching accuracy (Acc.), defined as the percentage of fine matchings with errors less than 5 pixels. Results in Table X demonstrate that the integration of learnable transformation alignment and the soft-matching approach enhances matching accuracy on the SemanticKITTI dataset. Consequently, our method facilitates the network’s comprehensive understanding of the relationship between image and point cloud data, learning effective representation to process information from each domain.
@@ IV-G Accuracy of Image-to-Point Cloud Calibration
We compared our approach with other leading methods for image-to-point cloud registration/calibration [58, 49, 60]. Following standard practices in this field, we used the SemanticKITTI dataset to measure the accuracy of our method, employing metrics such as Relative Translational Error (RTE) and Relative Rotational Error (RRE) [49]. We provided a detailed analysis by illustrating the registration recall at various RTE and RRE thresholds on the SemanticKITTI dataset, as shown in Fig. 9. The results clearly indicate the enhanced performance of our technique in terms of registration accuracy.

### v4

@@ IV-D Panoptic Segmentation
To assess the instance-level features learned by our method, we further fine-tuned our pre-trained models for the task of panoptic segmentation.
Network Architectures. In this part, we chose MINKUnet-34 [41] as the 3D point-wise feature backbone. For the fine-tuning process, semantic and instance heads are integrated with the pre-trained 3D backbone, followed by clustering post-processing to identify the instances [10]. This clustering stage leverages semantic predictions to exclude background points, focusing solely on foreground instances such as vehicles, pedestrians, and cyclists. Once the irrelevant points are removed, the remaining points undergo a clustering process to discern distinct instances based on the features from the instance head. For the clustering phase, we opt for the mean shift algorithm [60] with a set bandwidth of 1.2 and a minimum cluster size requirement of 50 points. The other architectures and settings are the same as Section IV-B during pre-training.
Protocol for Fine-tuning on Downstream Tasks. Our methodology employed the AdamW optimizer, utilizing a max learning rate set at 0.2. In the context of panoptic segmentation applied to the SemanticKITTI dataset, the fine-tuning of our model was conducted across various subsets of the annotated training data, specifically 0.1%, 1%, 10%, 50%, and the full 100% subset. For the nuScenes dataset, the fine-tuning process used both the complete training set and the mini training subset that is provided. The performance of the fine-tuned model was then assessed on the complete validation set of all datasets. We fine-tuned the model for fixed 50 epochs.
Results of Fine-tuning. We pre-trained the backbone on the SemanticKITTI dataset and fine-tuned it on different percentage subsets of the SemanticKITTI dataset. As shown in Table VI, our method is consistently better than previous self-supervised pre-training approaches. When the segmentation model is trained with fewer labels, we can observe an obvious improvement compared to the baseline of scratch training.
Generalization of Learned Features. Our study also investigated the adaptability of learned features focusing on panoptic segmentation tasks. We initially pre-trained the 3D backbone on the SemantiKITTI dataset and subsequently fine-tuned it on both the full and mini training sets of nuScenes. As presented in Table VII, while all approaches enhance performance on the nuScenes dataset, the features generated through our technique demonstrate greater adaptability when employed in a different domain dataset. Furthermore, in comparison to supervised pre-training on SemanticKITTI, our approach exhibits improved results. This finding underscores the enhanced effectiveness and potential of our method in scenarios of transfer learning, surpassing the conventional supervised pre-training methods.
@@ IV-E Ablation Studies
In our ablation studies on the nuScenes dataset, we adhered to the evaluation protocol from ALSO [15]. This involves splitting the nuScenes training set into two parts: ablation-train and ablation-val. Fine-tuning was performed with just 1% of annotations from the ablation-train set. This ensures parameter tuning does not involve the validation set, reserved for comparing against other methods. We limited the training period for these studies to 100 epochs.
@@
Effect of Key Components. We analyze the impact of each component incorporated into our method in Table VIII. We can observe that our proposed key components benefit the pre-training framework in learning useful representations and yield better performance in downstream tasks. For example, the introduction of learnable transformation alignment increases the mIoU from 35.90% to 36.52%. This indicates that allowing the model to adaptively transform features enhances its ability to capture the inherently non-linear and complex relationship between point cloud data and RGB images. Besides, the supervision of advanced tasks, including overlapping area detection and pose estimation also prompts the network to fully understand the two modalities and learn useful representation. Then we analyze the effect of learnable transformation alignment in detail.
w/o LTA | w/ LTA | |
Similarity | 0.55 0.17 | 0.69 0.13 |
@@
We also provide a quantitative comparison of similarities between the features of two modalities in Table IX. Specifically, we computed the cosine similarity between the features of the points and corresponding pixels with and without linear feature transformation, respectively. We can observe that directly computing the cosine similarities yields smaller values for the corresponding points and pixels. However, applying the learnable feature transformation bridges the gap between the two modalities and brings about more reasonable similarities.
@@ IV-F Feature Matching Accuracy
We present a comparative analysis of feature matching accuracy using various feature alignment and point-to-pixel matching methods. The mean and standard deviation of error distances are detailed in Table X. To assess the precision of different methods, we calculated the matching accuracy (Acc.), defined as the percentage of fine matchings with errors less than 5 pixels. The results in Table X demonstrate that the integration of learnable transformation alignment and the soft-matching approach enhances matching accuracy on the SemanticKITTI dataset. Consequently, our method facilitates the network’s comprehensive understanding of the relationship between image and point cloud data, learning effective representation to process information from each domain.
@@ IV-G Accuracy of Image-to-Point Cloud Calibration
We also compare our approach with other leading methods for image-to-point cloud registration/calibration [62, 53, 64]. Following standard practices in this field, we used the SemanticKITTI dataset to measure the accuracy of our method, employing metrics such as Relative Translational Error (RTE) and Relative Rotational Error (RRE) [53]. We provide a detailed analysis by illustrating the registration recall at various RTE and RRE thresholds on the SemanticKITTI dataset, as shown in Fig. 9. The results clearly indicate the enhanced performance of our technique in terms of registration accuracy.